OpenStudy (anonymous):

write (7-i)/(3-2i) in standard form for a complex number

OpenStudy (anonymous):

do you know what the conjugate is?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

(x+y) and (x-y) are conjugates. they're the same except for the sign. the deal with conjugates is that when you multiply them, you don't get a middle term. (x+y)(x-y) = x^2 + xy - xy - y^2 = x^2 - y^2 to divide by a complex number, you simply multiply the top and bottom by the conjugate of the bottom. this get's rid of the i and you're left with a real number on the bottom.

OpenStudy (anonymous):

so how do i begin to solve this?

OpenStudy (anonymous):

what number are you dividing by?

OpenStudy (anonymous):

what is its conjugate?

OpenStudy (anonymous):

multiply top and bottom by the conjugate. give it a shot and if you get stuck, i'm here.

OpenStudy (anonymous):

\[\frac{ 7-i }{ 3-2i }\] so that is how it is set up...now what? lol sorry !

OpenStudy (anonymous):

what is on the bottom? what is its conjugate?

OpenStudy (anonymous):

the 3-2i?

OpenStudy (anonymous):

yes... now what is its conjugate?

OpenStudy (anonymous):

so multiply the top and bottom by the 3-2i

OpenStudy (anonymous):

no... by the conjugate of that.

OpenStudy (anonymous):

x-y?

OpenStudy (anonymous):

change the sign inbetween... that's all you do!

OpenStudy (anonymous):

so the conjugate of 3 - 2i is just: 3 change the sign 2i

OpenStudy (anonymous):

what should the sign be?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

yeah!!! so multiply top and bottom by 3 + 2i...

OpenStudy (anonymous):

\[\frac{ 7-i }{ 3-2i }\cdot \frac{ 3+2i}{ 3+2i }=\frac{ (7-i)(3+2i)}{ 3^{2}+2^{2} }\] you can finish it off...

OpenStudy (anonymous):

so i got.. \[\frac{ 21-2i ^{2} }{ 9-4i ^{2}}\] ??

OpenStudy (anonymous):

almost... \[\frac{ (7-i)(3+2i) }{ 3^2-4i^2 }=\frac{ (21-2i^2+14i-3i) }{ 9-4(-1) }=\frac{ (21-2(-1)+14i-3i) }{ 9+4 }\] can you get it from here?

OpenStudy (anonymous):

\[\frac{ 12 }{ 13 }\] ?

OpenStudy (anonymous):

\[\frac{ (23+11i) }{ 13 }=\frac{ 23 }{ 13 }+\frac{11}{ 13 }i\] sorry... got to run! good luck!

OpenStudy (anonymous):

ok thanks!!